OpenStudy (anonymous):

A block is sliding down a inclined plane that has the angle 30° and the coefficient of sliding friction is 0,2 what is the acceleration of the block?

OpenStudy (anonymous):

this one is a little bit more complicated than your previous problem do you understand what the coefficient of friction tells you ?

OpenStudy (anonymous):

no not exactly

OpenStudy (anonymous):

well that is essential to solve this problem

OpenStudy (anonymous):

ok what i've got is: fk = N*muk = 9,8cos30 * 0,2 = 1,54 .. but i cant use a=F/m because i dont have the mass.